[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Oct 8 03:04:24 UTC 2017

HSU Shift Log for Sat Oct 07 2017 23:04:19 GMT-0400 (EDT)

Shift: R

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Bonefish froze and X-server had to be restarted. Satellite  
loops continued to load slowly after that.

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171008/a2540f80/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list